OpenStudy (credmond):

convert the equation into polar form. y=3x^2

OpenStudy (anonymous):

What is the conversion formula for x in terms of r and theta and the conversion formula for y in terms of r and theta?

OpenStudy (credmond):

x= (r)cos(theta) y= (r)sin(theta)

OpenStudy (anonymous):

Exactly ... now just substitute. those in and simplify

OpenStudy (credmond):

I don't know how to set it up :(

OpenStudy (anonymous):

\[r \sin \theta = 3r^2 \cos^2 \theta\] is probably sufficient, although it can be simplified further in the form r = f(theta)

OpenStudy (anonymous):

One could end up with polar equation r = (1/3)tanθsecθ